Combinatorial Optimization

K. Subramani, LCSEE, West Virginia University February 15, 2013

1. Bipartite graphs and Vertex covers. K_{3}.

2. Konig’s theorem. Compare with Konig’s orginal proof.

3. Hall’s theorem.

4. The consecutive one’s property.